# SEED-X: Multimodal Models with Unified Multi-granularity Comprehension and Generation

### v1

@@ Abstract
The rapid evolution of multimodal foundation model has demonstrated significant progresses in vision-language understanding and generation, e.g., our previous work SEED-LLaMA. However, there remains a gap between its capability and the real-world applicability, primarily due to the model’s limited capacity to effectively respond to various user instructions and interact with diverse visual data. In this work, we focus on bridging this gap through integrating two enhanced features: (1) comprehending images of arbitrary sizes and ratios, and (2) enabling multi-granularity image generation. We present a unified and versatile foundation model, namely, SEED-X, which is able to model multi-granularity visual semantics for comprehension and generation tasks. Besides the competitive results on public benchmarks, SEED-X demonstrates its effectiveness in handling real-world applications across various domains after instruction tuning. We hope that our work will inspire future research into what can be achieved by versatile multimodal foundation models in real-world applications. The models, codes, and datasets will be released in https://github.com/AILab-CVC/SEED-X.
@@ 1 Introduction
In this paper, we introduce SEED-X, a unified and versatile multimodal foundation model as a follow-up work of SEED-LLaMA, which seamlessly integrates the features mentioned above. After different instruction tuning, SEED-X can function as various multimodal AI assistants in the real world, capable of addressing various user needs through generating proper texts and images as shown in Fig. 1. Specifically, our instruction-tuned models can act as an interactive designer, generating images while illustrating creative intent, offering modification suggestions and showcasing visualizations based on user’s input images. Additionally, they can act as knowledgeable personal assistants, comprehending images of various sizes and providing relevant suggestions. Moreover, they can generate more diverse outputs, such as slide layouts for slide creation, and interleaved image-text content for storytelling. SEED-X signifies a notable advancement towards a versatile agent for users in the real world.
@@
To endow SEED-X with the aforementioned characteristics, our approach incorporates (1) a visual tokenizer to unify image comprehension and generation, where its multi-granularity de-tokenization phase facilitates image generation and high-precision image manipulation, and (2) an MLLM with dynamic resolution image encoding to enable the comprehension of images with arbitrary sizes and aspect ratios. Specifically, we utilize a pre-trained ViT as the visual tokenizer and train a visual de-tokenizer initialized from SDXL [20] to decode realistic images by taking the ViT features as input. Given the high-level ViT features as input, our visual de-tokenizer can reconstruct images semantically aligned with the original images. To realize the retention of fine-grained details of the input image to satisfy image manipulation, we further fine-tune the visual de-tokenizer to take an extra condition image as input in the latent space (See Fig. 2). The ViT features serve as a bridge to decouple the training of the visual (de-)tokenizer and the MLLM in SEED-X. The dynamic resolution image encoding divides an input image into sub-images and adds extrapolatable 2D positional embeddings to the ViT features of each sub-image, allowing the MLLM to scale to any image resolution. As for the image generation training objective, a fixed number of learnable queries are fed into the MLLM, where the output hidden states are trained to reconstruct the ViT features of the target images. During inference, the image de-tokenizer can take both the output features from the MLLM and the condition image provided by users as input, ensuring that the decoded image can possess high-level semantics that meet the multimodal instructions and retain the low-level details.
We pre-train SEED-X on massive multimodal data, including image-caption pairs, grounded image-text data, interleaved image-text data, OCR data, and pure texts. We further apply multimodal instruction tuning to align SEED-X with human instructions across various domains, utilizing both existing datasets and newly collected datasets that cover image editing, text-rich, grounded and referencing QA, and slide generation tasks. The extensive evaluations on MLLM benchmarks demonstrate that our instruction-tuned model not only achieves competitive performance in multimodal comprehension, but also achieves state-of-the-art results in image generation compared to existing MLLMs on SEED-Bench-2 [21].
All the models, codes, and datasets will be made publicly available. We hope our work can bring insights to the community about the potential of multimodal foundation models in real-world scenarios through unifying multi-granularity comprehension and generation.
@@ 2 Related Work
With the rapid development of Multimodal Large Language Models (MLLM), recent studies have been working on unified MLLMs that are capable of multimodal comprehension and generation as shown in Tab. 1. Some work [15, 14, 13, 22, 23] perform multimodal autoregression with a unified next-word-prediction objective through using a discrete visual tokenizer. For example, SEED-LLaMA [15] pre-trains a discrete image tokenizer, where the visual codes can be decoded to realistic images using the off-the-shelf SD model, and performs multimodal pre-training and instruction tuning through predicting the next discrete token. Some research efforts [12, 24, 19] have delved into multimodal autoregression with continuous representations, where each image is tokenized into embeddings via a visual encoder, and then interleaved with text tokens for autoregressive modeling. During inference, the regressed visual embeddings will be decoded into an image by a visual decoder. Additionally, some studies [17, 16] enable image generation in a non-autoregressive manner through utilizing learnable queries to obtain visual representations from MLLMs, which are further fed into a image decoder to generate images. The latest work, Mini-Gemini, generates text prompts using MLLMs and then leverages the existing SDXL [20] to generate images.
Table 1: MLLMs that unify comprehension and generation listed by publication date. “Decoder Input” denotes the inputs for image generation, where “Features” means continuous features, “Token” represents discrete tokens, and “Text” implies text prompts.
@@
Emu 07/2023 Feature $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
CM3Leon 07/ 2023 Token $\times$ $\times$ $\times$ $\times$ $\times$
SEED-OPT 07/ 2023 Token $\times$ $\times$ $\times$ $\times$ $\times$
LaVIT 09/2023 Token $\times$ $\times$ $\times$ $\times$ $\checkmark$
NExT-GPT 09/2023 Feature $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
DreamLLM 09/2023 Feature $\checkmark$ $\times$ $\times$ $\times$ $\times$
SEED-LLaMA 10/2023 Token $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
VL-GPT 12/2023 Feature $\checkmark$ $\times$ $\times$ $\times$ $\times$
Gemini 12/2023 Token $\checkmark$ $\times$ - $\times$ $\times$
Emu2 12/2023 Feature $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
Unified-IO 2 12/2023 Token $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$
Mini-Gemini 03/2024 Text $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
@@
In this work, we present SEED-X, a unified and versatile foundation model, through upgrading SEED-LLaMA with enhanced capabilities for real-world applications. As shown in Tab. 1, SEED-X supports object detection and dynamic resolution image encoding for multi-granularity comprehension, and high-level instructional image generation and low-level image manipulation for multi-granularity generation, which are not realized in previous work.
@@ 3.3.1 Training Stage I: Multimodal Pre-training
As shown in Fig. 4, SEED-X adopts next-word prediction and image feature regression training objectives on interleaved visual and textual data. Specifically, we perform dynamic resolution encoding of each image in the multimodal sequence, and their features along with text tokens are fed into the pretrained LLM. In order to equip the model with detection and referencing abilities, we add 224 bbox tokens, designated for representing bounding box coordinates, represented by ‘<box_start> <loc-x_center> <loc-y_center> <loc-width> <loc-height> <box_end>’ with specical tokens at the beginning and end of the bounding box. The text and added bbox tokens are training through predicting the next token with cross-entropy loss.
@@ 4 Experiments
Table 2: Comparison for multimodal comprehension and generation on MLLM benchmarks. “Image Gen” denotes whether the model can generate images besides texts. “Dynamic Res” denotes whether images with dynamic resolutions are encoded for training and inference. “Single”, “Multi” and “Interleaved” denote evaluating the comprehension of single-image, multi-image, and interleaved image-text respectively. “Gen” denotes evaluating the generation of images, and “/” denotes the model’s inability to perform such evaluation. The best results are bold and the second are underlined.
Size LLM Image Gen Dynamic Res MMB SEED-Bench-2 MME
Single P1 P2 P3
Perce-
ption
Cog-
nition
Single Multi
Inter-
leaved
Gen Single Single
GPT-4v - GPT-4 $\times$ - 77.0 69.8 73.1 37.9 / 1409 517
Gemini Pro - - $\checkmark$ - 73.6 62.5 - - - 1609 540
Qwen-VL-Chat 10B Qwen-7B $\times$ $\times$ 61.8 50.3 37.4 38.5 / 1488 361
Next-GPT 13B Vicuna-7B $\checkmark$ $\times$ - 31.0 27.8 40.3 42.8 - -
Emu 14B LLaMA-13B $\checkmark$ $\times$ - 46.4 31.2 45.6 45.7 - -
SEED-LLaMA-I 14B LLaMA-13B $\checkmark$ $\times$ - 49.9 32.4 48.0 50.6 - -
LLaVA-1.5 8B Vicuna-7B $\times$ $\times$ 66.5 58.3 39.2 34.4 / 1506 302
XComposer-VL 8B InternLM-7B $\times$ $\times$ 74.4 66.5 50.0 29.0 / 1528 391
SPHINX-1k - LLaMA-13B $\times$ $\times$ 67.1 68.5 37.7 32.5 / 1560 310
Emu2-Chat 37B LLaMA-33B $\times$ $\times$ 63.6 - - - / 1345 333
SEED-X-I 17B LLaMA-13B $\checkmark$ $\times$ 77.8 66.8 57.1 40.5 61.6 1520 338
SEED-X-I 17B LLaMA-13B $\checkmark$ $\checkmark$ 75.4 64.2 57.3 39.8 62.8 1457 321
@@ 4.1 Quantitative Evaluation
We evaluate SEED-X-I on benchmarks specifically designed for evaluating MLLMs, since recent work [31, 21] point out that traditional VQA benchmarks are not tailored for evaluating MLLMs with open-from output. As shown in Tab. 2, SEED-X-I achieves competitive performance in both the image comprehension and generation tasks. For example, SEED-X-I with dynamic resolution inputs achieves results on MMBench that are second only to GPT-4V, while SEED-X-I with fixed image size surpasses the performance of GPT-4V. Since MLLM Benchmarks lack a large number of data with unusual aspect ratios, the advantages of dynamic resolution encoding cannot be fully demonstrated, resulting in SEED-X-I with fixed image size achieving overall better performance. SEED-X-I also shows promising results for comprehending multi-image and interleaved image-text content in SEED-Bench-2 [32]. Compared with previous work [18, 15, 16] that unify comprehension and generation within an LLM, SEED-X-I achieves the state-of-the-art performance in P3 level of SEED-Bench-2 including the evaluation of text-to-image generation, next image prediction and text-image creation.
@@ Appendix A Pre-training and Instruction Tuning Datasets
LAION-COCO [35] (Re-caption), SAM [36] (Re-caption),
LAION-Aesthetics[26], Unsplash [27], JourneyDB [37], CapFusion [38],
@@
LAION-COCO [35] (Re-caption), SAM [36] (Re-caption),
LAION-Aesthetics[26], Unsplash [27], JourneyDB [37]
@@
Instructpix2pix [29], MagicBrush [30],
Openimages [56]-editing (In-house), Unsplash [27]-editing (In-house)

### v2

@@ Abstract
The rapid evolution of multimodal foundation model has demonstrated significant progresses in vision-language understanding and generation, e.g., our previous work SEED-LLaMA. However, there remains a gap between its capability and the real-world applicability, primarily due to the model’s limited capacity to effectively respond to various user instructions and interact with diverse visual data. In this work, we focus on bridging this gap through integrating two enhanced features: (1) comprehending images of arbitrary sizes and ratios, and (2) enabling multi-granularity image generation. We present a unified and versatile foundation model, namely, SEED-X, which is able to model multi-granularity visual semantics for comprehension and generation tasks. Besides the competitive results on public benchmarks, SEED-X demonstrates its effectiveness in handling real-world applications across various domains after instruction tuning. We hope that our work will inspire future research into what can be achieved by versatile multimodal foundation models in real-world applications. The models, codes, and datasets are released in https://github.com/AILab-CVC/SEED-X [^1] This is the v2 version. We added benchmark results (without updating models) and ablation study..
@@ 1 Introduction
In this paper, we introduce SEED-X, a unified and versatile multimodal foundation model as a follow-up work of SEED-LLaMA, which seamlessly integrates the features mentioned above. It is important to emphasize that integrating all these characteristics into a single foundation model is by no means trivial, as shown in Table 1, since none of the previous works support all of these features.
After different instruction tuning, SEED-X can function as various multimodal AI assistants in the real world, capable of addressing various user needs through generating proper texts and images as shown in Fig. 1. Specifically, our instruction-tuned models can act as an interactive designer, generating images while illustrating creative intent, offering modification suggestions and showcasing visualizations based on user’s input images. Additionally, they can act as knowledgeable personal assistants, comprehending images of various sizes and providing relevant suggestions. Moreover, they can generate more diverse outputs, such as slide layouts for slide creation, and interleaved image-text content for storytelling. SEED-X signifies a notable advancement towards a versatile agent for users in the real world.
@@
To endow SEED-X with the aforementioned characteristics, our approach incorporates (1) a visual tokenizer to unify image comprehension and generation, where its multi-granularity de-tokenization phase facilitates image generation and high-precision image manipulation, and (2) an MLLM with dynamic resolution image encoding to enable the comprehension of images with arbitrary sizes and aspect ratios. Specifically, we utilize a pre-trained ViT as the visual tokenizer and train a visual de-tokenizer to decode realistic images by taking the ViT features as input. To realize the retention of fine-grained details of the input image to satisfy image manipulation, we further fine-tune the visual de-tokenizer to take an extra condition image as input in the latent space (See Fig. 2). The ViT features serve as a bridge to decouple the training of the visual (de-)tokenizer and the MLLM. The dynamic resolution image encoding divides an input image into sub-images and adds extrapolatable 2D positional embeddings to the ViT features of each sub-image, allowing the MLLM to scale to any image resolution. For image generation, a fixed number of learnable queries are fed into the MLLM, where the output hidden states are trained to reconstruct the ViT features of the target images. During inference, the image de-tokenizer can take both the output features from the MLLM and the condition image provided by users as input, ensuring that the decoded image can possess high-level semantics that meet the multimodal instructions and retain the low-level details.
We pre-train SEED-X on massive multimodal data, including image-caption pairs, grounded image-text data, interleaved image-text data, OCR data, and pure texts. We further apply multimodal instruction tuning to align SEED-X with human instructions across various domains, utilizing both existing datasets and newly collected datasets that cover image editing, text-rich, grounded and referencing QA, and slide generation tasks. The extensive evaluations on MLLM benchmarks demonstrate that our instruction-tuned model not only achieves competitive performance in multimodal comprehension, but also exhibits excellent instruction-following capabilities for image generation.
All the models, codes, and datasets are made publicly available. We hope our work can bring insights to the community about the potential of multimodal foundation models in real-world scenarios through unifying multi-granularity comprehension and generation.
@@ 2 Related Work
With the rapid development of Multimodal Large Language Models (MLLM), recent studies have been working on unified MLLMs that are capable of multimodal comprehension and generation as shown in Tab. 1. Some work [15, 14, 13, 20, 21, 22, 23, 24] utilize a discrete visual tokenizer to perform multimodal autoregression with a unified next-word-prediction objective or masked visual token prediction. Some research efforts [12, 25, 19] have delved into multimodal autoregression with continuous representations, where each image in the multimodal sequence is tokenized into embeddings via a visual encoder, and then interleaved with text tokens for autoregressive modeling. During inference, the regressed visual embeddings will be decoded into an image by a visual decoder. Additionally, some studies [17, 16] enable image generation in a non-autoregressive manner through utilizing learnable queries to obtain visual representations from MLLMs, which are further fed into a image decoder to generate images. Mini-Gemini, generates text prompts using MLLMs and then leverages the existing SDXL [26] to output images.
Although these work have achieved competitive results on various academic benchmarks, such as VQA and text-to-image generation, the accuracy and diversity of their generated content still fall short of real-world needs, since they do not meet the requirements of modeling multi-granularity visual semantics for comprehension and generation task. As shown in Tab. 1, we identify several significant characteristics essential for real-world applications including object detection and dynamic resolution image encoding for multi-granularity comprehension, as well as high-level instructional image generation and low-level image manipulation for multi-granularity image generation. Notably, none of the previous works fully support all of these characteristics. In this work, we present SEED-X, a unified and versatile foundation model, which effectively incorporate the aforementioned characteristics for real-world applications.
Table 1: MLLMs that unify comprehension and generation and whether they support significant characteristics essential for real-world applications. “Decoder Input” denotes the inputs for image generation, where “Features” means continuous features, “Token” represents discrete tokens, “Text” implies text prompts.
@@
Emu 07/2023 Feature $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
CM3Leon 07/ 2023 Token $\times$ $\times$ $\checkmark$ $\times$ $\times$
SEED-OPT 07/ 2023 Token $\times$ $\times$ $\checkmark$ $\times$ $\times$
LaVIT 09/2023 Token $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
NExT-GPT 09/2023 Feature $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
DreamLLM 09/2023 Feature $\times$ $\times$ $\checkmark$ $\times$ $\times$
SEED-LLaMA 10/2023 Token $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
VL-GPT 12/2023 Feature $\times$ $\times$ $\checkmark$ $\times$ $\times$
Gemini 12/2023 Token $\times$ - $\checkmark$ $\times$ $\times$
Emu2 12/2023 Feature $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
Unified-IO 2 12/2023 Token $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$
Mini-Gemini 03/2024 Text $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
@@ 3.3.1 Training Stage I: Multimodal Pre-training
As shown in Fig. 4, SEED-X adopts next-word prediction and image feature regression training objectives on interleaved visual and textual data. Specifically, we perform dynamic resolution encoding of each image in the multimodal sequence, and their features along with text tokens are fed into the pretrained LLM. In order to equip the model with detection and referencing abilities, we add 224 bbox tokens, designated for representing bounding box coordinates, represented by <box_start> <loc-x_center> <loc-y_center> <loc-width> <loc-height> <box_end> with special tokens at the beginning and end of the bounding box. The text and added bbox tokens are trained through predicting the next token with cross-entropy loss.
@@
Table 2: Comparison on multimodal understanding benchmarks. “Und.” and “Gen.” denote “understanding” and “generation”, respectively.
[…]
Und. Only LLaVA-v1.51.51.51.5-Phi-$1.5$ [23] $84.1$ $1128.0$ $75.3$ $56.5$ $30.7$
MobileVLM [33] $84.5$ $1196.2$ $53.2$ $56.1$
MobileVLM-V2 [34] $84.3$ $1302.8$ $57.7$ $59.3$
LLaVA-Phi [35] $85.0$ $1335.1$ $59.8$ $71.4$ $28.9$
LLaVA [36] $76.3$ $809.6$ $38.7$ $33.5$ $25.5$
LLaVA-v1.51.51.51.5 [37] $\mathbf{85.9}$ $\mathbf{1510.7}$ $64.3$ $58.6$ $\mathbf{78.5}$ $\mathbf{62.0}$ $35.4$ $31.1$
InstructBLIP [38] - - $36.0$ $53.4$ $49.2$ $26.2$
IDEFICS-$9$[39] - - $48.2$ $50.9$ $38.4$
Qwen-VL-Chat [5] - $1487.5$ $60.6$ $58.2$ $78.2$ $57.5$
Und. and Gen. DreamLLM [17] - - - - $72.9$ $36.6$
LaVIT [20] - - - - $66.0$ $46.8$
Emu [18] - - - - $52.0$
NExT-GPT [40] - - - - $66.7$
Gemini-Nano-1 [41] - - - - $62.7$ $26.3$
LWM [42] $75.2$ $55.8$ $44.8$ $9.6$
SEED-X $84.1$ $1457.0$ $\mathbf{70.1}$ $\mathbf{66.5}$ $71.2$ $49.1$ $\mathbf{35.6}$ $\mathbf{43.0}$
@@ 4.1 Quantitative Evaluation
Multimodal Comprehension. We evaluate the multimodal comprehension capabilities of SEED-X-I on widely recognized image-based vision-language benchmarks, which include VQAv2 [43], GQA [44], POPE [45], MME [46], SEED [47], MMB [48], MM-Vet [49], and MMMU [50]. As listed in Tab. 2, SEED-X-I achieves competitive performance across various benchmarks, even when compared to MLLMs specifically designed for multimodal comprehension.
Image Generation. We evaluate the image generation capabilities of SEED-X-I on GenEval [51], which is a challenging benchmark to evaluate compositional image properties such as object co-occurrence, position, count, and color. As shown in Tab. 3, SEED-X obtains 51% overall accuracy, demonstrating the model’s excellent instruction-following capabilities for image generation.
Table 3: Evaluation of text-to-image generation ability on GenEval benchmark. “Und.” and “Gen.” denote “understanding” and “generation”, respectively.
[…]
Gen. Only LDM [52] $0.92$ $0.29$ $0.23$ $0.70$ $0.02$ $0.05$ $0.37$
SDv1.51.51.51.5 [52] $0.97$ $0.38$ $0.35$ $0.76$ $0.04$ $0.06$ $0.43$
PixArt-$\alpha$ [53] $0.98$ $0.50$ $0.44$ $0.80$ $0.08$ $0.07$ $0.48$
SDv2.12.12.12.1 [52] $0.98$ $0.51$ $0.44$ $0.85$ $0.07$ $0.17$ $0.50$
DALL-E $2$ [54] $0.94$ $0.66$ $0.49$ $0.77$ $0.10$ $0.19$ $0.52$
SDXL [26] $0.98$ $0.74$ $0.39$ $0.85$ $0.15$ $0.23$ $0.55$
Und. and Gen. LWM [42] $0.93$ $0.41$ $0.46$ $0.79$ $0.09$ $0.15$ $0.47$
SEED-X $0.96$ $0.65$ $0.31$ $0.80$ $0.18$ $0.14$ $0.51$
@@ 4.2 Qualitative Evaluation
Figure 7: Ablation study on the number of visual tokens and trainable parameters for training visual de-tokenizer.
4.3 Ablation Study
In this section, we perform ablation studies on the training of our visual de-tokenizer and the pre-training of SEED-X to enable a MLLM for image generation.
For visual de-tokenization, N visual embeddings (after average pooling) from the ViT tokenizer are fed into a learnable module as the inputs of the U-Net of the pre-trained SD-XL. We perform an ablation study on the number of visual tokens and the learnable parameters of the SD-XL U-Net, where keys and values within the U-Net are optimized if not specified with “fully fine-tune”. As shown in Fig. 7, we can observe that more visual tokens can result in better reconstruction of the original images. For example, the decoded images from 256 visual embeddings can recover the characters’ postures of the original images, while decoded images from 32 visual embeddings have already lost the original structure of the scene. We further observe that fully fine-tuning the parameters of the SD-XL U-Net can lead to distortions in image details, such as the woman’s feet, compared to only training the keys and values within the U-Net. In SEED-X, we use N = 64 visual embeddings to train the visual de-tokenizer and only optimize the keys and values within the U-Net (See below for an explanation of why we do not choose N = 256).
To enable MLLM for image generation, we employ N learnable queries to obtain the output visual representations from the LLM, which are trained to reconstruct N visual embeddings from the ViT tokenizer with a learnable module. We first perform an ablation study on the number of learnable queries. The images generated by the MLLM based on the input caption are shown in Fig. 8. We can observe that using 256 learnable queries to reconstruct 256 visual embeddings can lead to distortion in the generated images compared with N = 64. This occurs because regressing more visual features is more challenging for the model, even though 256 visual embeddings from the de-tokenizer can better reconstruct images, as demonstrated in the previous ablation study. We also observe that, compared to learning a one-layer cross-attention for reconstructing image features, a multi-layer resampler (multi-layer cross-attention) yields less satisfactory performance, which can happen due to the lack of more direct regularizations on the hidden states of the LLM. We further optimize the visual de-tokenizer by using the reconstructed visual embeddings from the MLLM as input instead of ViT features, but the generated images exhibit a more monotonous appearance. It demonstrates the effectiveness of utilizing the ViT Tokenizer as the bridge to decouple the training of visual de-tokenizer and the MLLM for image generation.
Figure 8: Ablation study on the number of visual tokens, model architecture and optimization targets during pre-training SEED-X for image generation.
@@ Appendix A Pre-training and Instruction Tuning Datasets
LAION-COCO [56] (Re-caption), SAM [57] (Re-caption), Unsplash [29],
LAION-Aesthetics[28], JourneyDB [58], CapFusion [59],
@@
LAION-COCO [56] (Re-caption), SAM [57] (Re-caption), Unsplash [29],
LAION-Aesthetics[28], JourneyDB [58]
@@
Instructpix2pix [31], MagicBrush [32], Openimages [76]-editing (In-house),
Unsplash [29]-editing (In-house)